Looking at the data during the last month (starting from April 12th), there seem to be a hint of etalon behavior in both the arms. This behavior was not spotted in the past analysis (#45832) since the considered time period was shorter and the effect was overlapped by the TCS working point change (#45727).
Figure 1 and 2 show the power on B7 and B8, respectively, both normalized by the recycling cavity power (B4) as a function of the temperature as sensed by the RH thermometer. Although a complete etalon fringe has not been performed, however the peak to peak power difference results to be of the order of 1.2% for the North arm for a temperature variation of 0.15K, while for the West arm the power difference is of about 2.4% with the same temperature variation. These values are compatible with what was already found in the past (#39767). Notice that the temperature trend has been double checked with the frequency variation of the drum lines, calibrated using the same coefficients reported in #39767(see figures 3 and 4).
In figure 5 the Optical gain of the cavities is plotted versus the temperature of NI and WI. Although the OG variation cannot be attributed only to the etalon effect, there seems to be a trend identifing a region around (18.95, 19.05) as the (N,W) temperatures maximizing the OG. However, this optimal point can depend on the current ITF working point.
It could be that the detrimental effect on the optical gain starts when the arms go through a different fringe condition, giving rise to a finesse asymmetry.